OpenStudy (anonymous):

Please help, medals will be given!!! Solve x + 2y – z = 0 2x + 5y + z = 3 3x – y + z = 29 What is the value of x?

OpenStudy (amistre64):

we can use eq1 to eliminate the z parts, then it reduces to 2 eqs in x and y

OpenStudy (anonymous):

All you are doing is bringing everything to the other side of the = sine

OpenStudy (anonymous):

So the first one would be x=0+z-2y

OpenStudy (anonymous):

The answer choices are just single numbers.....

OpenStudy (amistre64):

bean seems to be suggesting a substituion method ... while mine is an elimination method

OpenStudy (anonymous):

Can you post the answer choices?

OpenStudy (amistre64):

the study part of this would be to address what thoughts maria may have on a solution process.

OpenStudy (anonymous):

Yes that sounds like a good idea I was going to help her solve the first one so she knew how to do it.

OpenStudy (anonymous):

Its a hidden message assignment but the answers left are: 28,14,-2,-4,5,-5,-3,-10,6,and 8

OpenStudy (anonymous):

If anyone goes to k12 its the unit 2 extracredit assignment for algebra 2, mrs. shuler....

OpenStudy (amistre64):

so what have you tried in working a solution?

OpenStudy (anonymous):

elimination i think

OpenStudy (amistre64):

elimination is a valid approach :) can you write up what your steps were? i can verify your attempts and correct them if need be

OpenStudy (anonymous):

I didn't know where to begin on this one, the last one I did was two equation not three.

OpenStudy (amistre64):

thats fine, one thing i notice is that the first equation has a -z while the others have a +z we can use the first equation to produce 2 new equations that have the z parts eliminated

OpenStudy (anonymous):

If you combine each equation like x+2y - z = 0 2x+5y+z = 3 _________________ Did you learn any thing like this?

OpenStudy (amistre64):

x + 2y – z = 0 2x + 5y + z = 3 --------------- 3x +7y +0 = 3 x + 2y – z = 0 3x – y + z = 29 ---------------- 4x +y + 0 = 29 now this should be something you are used to dealing with: 3x + 7y = 3 4x + y = 29

OpenStudy (anonymous):

so you do the first one seperatly with the other two??

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

would you agree that the new setup in xy, we can eliminate the y parts fairly easily? -7 times eq2 should do it

OpenStudy (anonymous):

then what

OpenStudy (amistre64):

then you eliminate the ys when you add the results .... that is the elimination process after all

OpenStudy (amistre64):

or if substitution is more your style, thats fine too .... the answer doesnt change

OpenStudy (anonymous):

so 8?

OpenStudy (amistre64):

let me check :)

OpenStudy (amistre64):

i get 8 as well

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

Thanks have a nice day!

OpenStudy (amistre64):

good luck ;)